OpenStudy (anonymous):

Which of the following equations best represents the trend line for the scatter plot shown below

OpenStudy (anonymous):

y = 1/3x + 8 y = 3x + 5 y = 3x + 8

OpenStudy (anonymous):

answers^

OpenStudy (anonymous):

Looks like y-intercept is at y=5, so it's y=3x+5